# King’s College London, U.K

### 11 Mar 2024 v1

@@ Abstract
We demonstrate that Assembly Theory, pathway complexity, the assembly index, and the assembly number are subsumed and constitute a weak version of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity reliant on an approximation method based upon statistical compression, their results obtained due to the use of methods strictly equivalent to the LZ family of compression algorithms used in compressing algorithms such as ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT’s assembly index and their use had already been reported in successful application to separating organic from non-organic molecules, and the study of selection and evolution. Here we exhibit and prove the connections and full equivalence of Assembly Theory to Shannon Entropy and statistical compression, and AT’s disconnection as a statistical approach from causality. We demonstrate that formulating a traditional statistically compressed description of molecules, or the theory underlying it, does not imply an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by selection and evolution, when lacking in logical consistency and empirical evidence. We argue that in their basic arguments, the authors of AT conflate how objects may assemble with causal directionality, and conclude that Assembly Theory does nothing to explain selection or evolution beyond known and previously established connections, some of which are reviewed here, based on sounder theory and better experimental evidence.
@@ 1 Introduction
Assembly Theory (AT) has recently garnered significant attention in some academic circles and in the scientific media. Responding to how open-ended forms can emerge from matter without a blueprint, AT purports to explain and quantify selection and evolution. The central claim of Assembly Theory advanced in [1, 2] is that objects with a high assembly index “are very unlikely to form abiotically”. While this has been contested in [3], we will demonstrate the connections of Assembly Theory and its assembly index to well-established theories that the authors claim are different and unrelated to Assembly Theory [4]. We will show that the claims advanced by its authors that Assembly Theory unifies life and biology with physics [2, 5] rely on a circular argument and the use of a popular compression algorithm, with no empirical or logical support to establish deeper connections to selection and evolution not already made via more meaningful experiments and based on more solid theoretical grounds [6, 7]. We demonstrate how Assembly Theory is a weak version of algorithmic complexity (a.k.a. Kolmogorov-Solomonoff-Chaitin complexity) [8, 9, 10, 11], denoted by $\mathbf{K}$, that its proposed measure is an LZ compression scheme and, unsurprisingly, asymptotically bounded by (Shannon) Entropy in a fashion that is in no way dissimilar to previous information and graph-theoretic approaches to chemical complexity. We conclude that Assembly Theory cannot offer a different or better explanation of selection or evolution than the connections already established.
2 Dissasembling Assembly Theory: Main Concerns
@@ 3 Conclusions
On the other hand, here we have proven that the main components and substance of Assembly Theory are based upon and subsumed into the theory of algorithmic complexity. Based on their high and low copy number feature driving pathway complexity and their assembly number, an object with low assembly index will have low Entropy and high LZ compressibility, and an object with high assembly index will have high Entropy and low LZ compressibility with respect to the size of the object in the same proportions and driven by precisely the same copy-counting mechanism (on which every other popular statistical compression is also based upon). This means that Assembly Theory is a weak approximation to algorithmic complexity. The latter has more discriminatory power because it can tell apart cases of high assembly index with low copy number, high Entropy and low LZ compressibility but with potentially causal compressible content [12] that can be described by short rewriting rules of the same type of but a non-trivial nature than those only taken into account by Assembly Theory. Thus, algorithmic complexity is more profoundly connected to causality [36, 37, 20, 38]. The claim that it can only be applied to strings is also false, as it has been applied to images [39], networks [7, 40, 41, 42], vectors, matrices, and tensors in multiple dimensions [43, 44, 45]; and even chemical structures before [17].

### 13 Mar 2024 v2

@@ Abstract
We demonstrate that Assembly Theory, pathway complexity, the assembly index, and the assembly number are subsumed and constitute a weak version of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity reliant on an approximation method based upon statistical compression, their results obtained due to the use of methods strictly equivalent to the LZ family of compression algorithms used in compressing algorithms such as ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT’s assembly index and their use had already been reported in successful application to separating organic from non-organic molecules, and the study of selection and evolution. Here we exhibit and prove the connections and full equivalence of Assembly Theory to Shannon Entropy and statistical compression, and AT’s disconnection as a statistical approach from causality. We demonstrate that formulating a traditional statistically compressed description of molecules, or the theory underlying it, does not imply an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by selection and evolution, when lacking in logical consistency and empirical evidence. We argue that in their basic arguments, the authors of AT conflate how objects may assemble with causal directionality, and conclude that Assembly Theory does not explain selection or evolution beyond known and previously established connections, some of which are reviewed here.
@@ 1 Introduction
Assembly Theory (AT) has recently garnered significant attention in some academic circles and in the scientific media. Responding to how open-ended forms can emerge from matter without a blueprint, AT purports to explain and quantify selection and evolution. The central claim of Assembly Theory advanced in [1, 2] is that objects with a high assembly index “are very unlikely to form abiotically”. While this has been contested in [3], we will demonstrate the connections of Assembly Theory and its assembly index to well-established theories that the authors claim are different and unrelated to Assembly Theory [4]. We will show that the claims advanced by its authors that Assembly Theory unifies life and biology with physics [2, 5] rely on a circular argument and the use of a popular compression algorithm, with no empirical or logical support to establish deeper connections to selection and evolution not already made before [6, 7]. We demonstrate how Assembly Theory is a weak version of algorithmic complexity (a.k.a. Kolmogorov-Solomonoff-Chaitin complexity) [8, 9, 10, 11], denoted by $\mathbf{K}$, that its proposed measure is an LZ compression scheme and asymptotically bounded by (Shannon) Entropy in a fashion that is in no way dissimilar to previous information and graph-theoretic approaches to chemical complexity. We conclude that Assembly Theory cannot offer a different or better explanation of selection or evolution than the connections already established.
2 Disassembling Assembly Theory: Main Concerns
@@ 3 Conclusions
On the other hand, here we have proven that the main components and substance of Assembly Theory are based upon and subsumed into the theory of algorithmic complexity. Based on their high and low copy number feature driving pathway complexity and their assembly number, an object with a low assembly index will have low Entropy and high LZ compressibility, and an object with a high assembly index will have high Entropy and low LZ compressibility with respect to the size of the object in the same proportions and driven by the same copy-counting mechanism (on which every other popular statistical compression is also based upon). This means that Assembly Theory is a weak approximation to algorithmic complexity. The latter has more discriminatory power because it can tell apart cases of high assembly index with low copy number, high Entropy, and low LZ compressibility but with potentially causal compressible content [12] that can be described by short rewriting rules of the same type of but a non-trivial nature than those only taken into account by Assembly Theory. Thus, algorithmic complexity is more profoundly connected to causality [36, 37, 20, 38]. The claim that it can only be applied to strings is also false, as it has been applied before to images [39], networks [7, 40, 41, 42], vectors, matrices, and tensors in multiple dimensions [43, 44, 45]; and even chemical structures [17].